OpenStudy (anonymous):

In an election 1200 votes are counted for 3 candidates. The second place candidate recieves 125 less votes than the winner and 200 more than the third place candidate. How many votes did the winner recieve

OpenStudy (anonymous):

So far I have done this: x = 2nd place votes 125+x = winner votes x-200 = third place votes But how do I go from these 3 expressions to 1 equation for how many votes the winner recieved?

OpenStudy (anonymous):

do 1200 divide by 3= 400 400+125=525 400-125=275

OpenStudy (anonymous):

hope this helped you